Fermi National Laboratory
 

Accelerator Update for April 27, 2007 - April 30, 2007

 

Friday April 27

The day shift began with Operations monitoring the Tevatron (TeV) store 5379 (80.76E30) and the Recycler stash (115.8E10), with the Antiproton Source (Pbar) stacking (47.8E10), with the Linac, Booster, and Main Injector (MI) operating normally, with NuMI and MiniBooNE taking beam, and with D-Zero (D0) and the Collider Detector at Fermilab (CDF) using the colliding beams.

At 9:30 AM, a Linac Klystron RF station (KRF3) tripped off. It reset.

At 9:51 AM, KRF4 tripped off. It wouldn’t reset. Experts investigated and had it up and running by 10:10 AM.

At 10:33 AM, all the power supplies at TeV sector F27 tripped off. This did not affect the colliding beams, only beam between MI and Pbar. Operations loaded a NuMI only timeline while experts investigated. LCW technicians reported that they had accidentally tripped off a water skid. Everything was back on by 11:38 AM.

A Pbar expert reported at 10:44 AM that he found a quadrupole power supply (D:Q724) that had no output. Operations contacted an expert.

MI experts began a study period at 10:50 AM.

A power supply expert had D:Q724 up and running by 10:55 AM.

At 12:07 PM, store 5379 suffered a quench at TeV sector C2 due to a HFU pre-fire indication.

Operations ramped the TeV at 1:57 PM.

Operations prepared CDF for an access at 3:04 PM.

Operations reported at 4:02 PM that the Recycler had stashed antiprotons three times during the day shift.

CDF came out of access at 4:21 PM.

Operations put the TeV into shot setup at 5:22 PM.

Operations established store 5382 at 7:17 PM with an initial luminosity of 192.1E30.

At 7:49 PM, Operations reported MI beam permit trips due to bad beam from Booster. Operators reset MI and tuned Booster.

At 10:07 PM, KRF4 tripped off. It reset.

Operations reported at 11:54 PM that the Recycler had stashed twice during the evening shift.

Saturday April 28

The midnight shift began with Operations monitoring store 5382 (100.38E30) and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 1:10 AM, KRF4 tripped off. It reset.

At 2:10 AM, MI tripped off due to bad beam positions. It reset.

At 2:21 AM, KRF3 tripped off. It reset.

At 2:51 AM, the Lithium Lens power supply tripped off due to a high temperature indication. It took 40 minutes before operators could reset the power supply.

At 5:04 AM, LRF3 tripped off. It reset.

At 6:27 AM, Operations reported problems with transfers of antiprotons from Pbar to the Recycler. Operators and Pbar experts investigated.

Operations reported at 7:54 AM that the Recycler had stashed three times during the midnight shift.

At 8:26 AM, all the MI RF stations tripped off due to a bad vacuum indication. Operators investigated, but found no problems. Everything reset.

At 10:41 AM, LRF3 tripped off. It reset.

At 11:25 AM, Operations reported a high temperature for the TeV sector A0 kicker room. Operations contacted the Duty Mechanic.

At 12:50 PM, Operations reported another Lithium Lens power supply trip due to high temperature. Operators contacted experts. The temperature alarm cleared and stacking resumed 41 minutes later.

Operations terminated store 5382 at 3 PM.

Operations put the TeV into shot setup at 3:12 PM.

Operations established store 5383 at 5:15 PM with an initial luminosity of 191.62E30. Operations reported that due to a problem with Electron Cooling some antiprotons were left in the Recycler.

At 6:59 PM, MIRF14 tripped off. It reset.

Operations reported at 8 PM that the Recycler had stashed three times during the last 12 hours. Accelerator status: Store 5383 was at 125.35E30, the stack was at 52.5E10, and the stash was at 93.5E10.

At 8:15 PM, two Recycler correction element power supplies tripped off (R:CE1010 & 102). There were no antiprotons lost. MI aborted at the same time. Operations contacted experts.

At 10:39 PM, the Lithium Lens power supply tripped off. Operators managed to reset the supply at 11:19 PM.

At 11:41 PM, LRF3 tripped off for the first of two times in 21 minutes. It reset.

At 11:44 PM, KRF4 tripped off. It reset.

Sunday April 29

The midnight shift began with Operations monitoring store 5383 and the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

At 1:09 AM, KRF5 tripped off. It reset.

At 2:49 AM, MI suffered the first of three trips in about one hour due to bad beam positions. It reset each time.

At 4:14 AM, the Lithium Lens power supply tripped off. Stacking resumed about 20 minutes later.

At 5:28 AM, KRF4 tripped off. It reset.

At 7:40 AM, LRF3 tripped off. It reset.

At 7:54 AM, Operations reported a Pbar Debuncher problem. Operators investigated and then contacted an expert.

Operations reported at 8 AM that the Recycler had stashed four times during the last 12 hours. Accelerator status: Store 5383 was at 52.32E30, the stack was at 25.2E10, and the stash was at 263.6E10.

A Pbar expert reported that there was a Debuncher problem that will need an access to fix, but since stacking was above 20 mA/hr the access could wait.

At 9:24 AM, KRF4 tripped off. It reset.

At 9:36 AM, the Lens power supply tripped off. Operators had it back on about 50 minutes later.

Operations reported at 10:33 AM that Booster problems were due to high temperatures in the low level RF room. The Duty electrician installed a portable AC unit.

At 12:08 PM, MIRF14 tripped off for the first of two times in six minutes. It reset.

At 12:30 PM, Operators installed a fan in the Booster LLRF room to direct cool air into the LLRF crate.

Operations terminated store 5383 at 2 PM and put the TeV into shot setup 15 minutes later.

At 2:45 PM, MIRF14 tripped off. It reset.

At 3:48 PM, a Recycler bulk power supply (C203) tripped off due to a LCW flow fault, but this wasn’t real. The supply was still operating normally. Operations contacted an expert and then reset the fault.

At 4:12 PM, Operations reported that it appeared that NuMI beam was falling out. However, it was Pbar slipstacked beam that was falling out. Operators discovered that the Booster damper had tripped off and caused the MI dampers to trip off, which caused all the other problems. A Booster expert came in and eventually everything reset okay.

Operations established store 5384 at 4:35 PM with an initial luminosity of 239.88E30.

Operations reported at 7:52 PM that the Recycler had stashed three times during the last 12 hours. Accelerator status: Store 5384 was at 136.09E30.

At 9:15 PM, Operations reported that MIRF14 had been having problems, was bypassed for a while because of the problems, brought back on line, but then bypassed again due to continuing problems.

At 9:19 PM, Operations reported that a Pbar Debuncher kicker problem (D:EKIK) had lowered the stack rate. An hour later Operations adjusted its timing and that solved the problem.

At 9:48 PM, LRF2 tripped off and would not reset due to a driver problem. Operators investigated and contacted an expert, but by then, about an hour later, the station had reset itself, tripped off again, and then came back up okay.

At 10:39 PM, Operations reported that the TeV aborted store 5384. Coincidentally, there was a loud rumble of thunder at the same time. Operators investigated and found that a lightning strike had caused the D0 abort to toggle.

Operations reported at 10:52 PM, that several other problems occurred during the lightning strike. D:EKIK developed a problem; operations called in an expert. The MI Power Supply Interlock Loop tripped off; it reset, but operators never discovered why it had tripped. Also, the Recycler indicated that it had suffered a quadrupole compensation interlock loop trip; it reset and no beam was lost.

Monday April 30

The midnight shift began with Operations monitoring the stash, with Pbar stacking, and with NuMI and MiniBooNE taking beam.

TeV experts began sextupole studies at 12:32 AM.

At 3:48 AM, KRF5 tripped off. It reset.

At 4:42 AM, the Lithium Lens power supply tripped off. Operations was able to reset it about an hour later.

At 7:03 AM, the NuMI beam permit tripped off due to bad beam positions. It reset.

Operations reported at 7:56 AM that a Pbar expert moved the target to a different spot.

Operations reported at 8:01 AM that the Recycler had stashed three times during the last 12 hours.

The Plans for Monday

The plans for today are to complete the TeV sextupole studies this morning and then establish a new store. The rest of the day will consist of stacking, stashing, and storing.


Accelerator Update Archive

More Information
For Tevatron luminosity charts and the current status of Fermilab's accelerators and detectors (live!), please go to Fermilab Now.

Comments and Suggestions
What do you think about the Accelerator Updates? Please send comments and suggestions to: accelupdates@fnal.gov.

last modified 04/30/2007    email Fermilab